OpenStudy (anonymous):

Use a repeating pattern to find the numeral that will be in the ones place of 2 (small upper numbers) 387?

OpenStudy (anonymous):

Consider the following: 2^1 = 2 2^2 = 4 2^3 = 8 2^4 = 16 2^5 = 32 2^6 = 64 2^7 = 128 2^8 = 256 2^9 = 512 2^10 = 1024 Look at the ones place in each number. You should see the following pattern: 2, 4, 8, 6, 2, 4, 8, 6, 2, 4... It simply repeats 2, 4, 8, 6 over and over. Whenever the exponent is a multiple of 4, the ones place is a 6. So, we just need to find where 387 is in this cycle. Using division to find the remainder of 387/4, we get 3. This means 384 is a multiple of 4, so 2^384 ends in a 6, and 387 is 3 greater than a multiple of 4, so 2^387 ends in an 8.